OpenStudy (lexilove13):

Help again! @jim_thompson5910

OpenStudy (lexilove13):

Solve the equation \[\frac{ 1 }{ 3x+9 } - \frac{ 2 }{ x+3 }=2\]

OpenStudy (lexilove13):

Answer choices: A. X=1 B. x=\[\frac{ -23 }{ 6 }\] C. x=\[\frac{ -5 }{ 2 }\] D. x=-5

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 1 }{ 3x+9 } - \frac{ 2 }{ x+3 }=2\] \[\large \frac{ 1 }{ 3(x+3) } - \frac{ 2 }{ x+3 }=2\] \[\large \frac{ 1 }{ 3(x+3) } - \frac{ 3*2 }{ 3(x+3) }=2\] \[\large \frac{ 1 }{ 3(x+3) } - \frac{ 6 }{ 3(x+3) }=2\] I'll let you finish

jimthompson5910 (jim_thompson5910):

you can't subtract unless the denominators are the same

OpenStudy (lexilove13):

OK. So it would be -5 or D! Thanks again. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

I'll finish up

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 1 }{ 3(x+3) } - \frac{ 6 }{ 3(x+3) }=2\] \[\large \frac{ 1 - 6 }{ 3(x+3) }=2\] \[\large \frac{ -5 }{ 3(x+3) }=2\] \[\large -5 =2*3(x+3)\] \[\large -5 =6x+18\] \[\large -5-18 =6x\] \[\large -23 =6x\] \[\large x = -\frac{23}{6}\]

OpenStudy (lexilove13):

That could have been bad! Thank you!

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

make sure you check any/all answers you get because sometimes with rational equations, possible solutions aren't really solutions at all

jimthompson5910 (jim_thompson5910):

in this case, x = -23/6 is actually a solution, but checking is always good